OpenStudy (anonymous):

Solve. Finding all solutions in (0,2π) cos(π/2 +x) + sin(x-π)=1

OpenStudy (maheshmeghwal9):

\[\cos (\frac{\pi}{2}+x)+\sin (x- \pi)=1.\]\[\implies - \sin x + [\sin- (\pi -x)]=1.\]\[\implies - \sin x- \sin (\pi -x)=1.\]\[\implies - \sin x -\sin x=1 \implies -2 \sin x=1 \implies \sin x=-\frac{1}{2}.\]

OpenStudy (maheshmeghwal9):

Now find 'x' in accordance with the interval \[\LARGE{\color{RED}{(0, 2 \pi).}}\]

OpenStudy (anonymous):

unfortunately i only have four anser choices i forgot to mention and they are either: a- 7π/6,11π/6 b- 7π/6 c-π/6, 7π/6 d- π/4, 3π/4

OpenStudy (anonymous):

I tell you how to do it just wait..

OpenStudy (maheshmeghwal9):

just figure out the answer by last equation which we got \[\implies \sin x=-\frac{1}{2}.\]

OpenStudy (maheshmeghwal9):

It would be option A

OpenStudy (maheshmeghwal9):

don't u get then wait for @waterineyes :D

OpenStudy (anonymous):

how did you get it from the original equation to −sinx+[sin−(π−x)]=1.

OpenStudy (anonymous):

is there a certain rule/identity you follow

OpenStudy (maheshmeghwal9):

I have written the solution above didn't u see it????????/

OpenStudy (maheshmeghwal9):

ya there are many identities but those which i used are as follows \[[1.] \space \space \cos (\frac{\pi}{2}+x)=- \sin x.\]\[[2.] \space \space \sin -( x) = - \sin x.\]\[[3.] \space \space \sin (\pi- x)=\sin x.\]

OpenStudy (anonymous):

ok, now i get how you got to it. For the first one it looks familiar, can you tell me what its called please.

OpenStudy (anonymous):

the rule/identity

OpenStudy (maheshmeghwal9):

Reduction formulae.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\(\cos (\frac{\pi}{2}+x)+\sin (x- \pi)=1\) \(\implies - \sin x + (\sin- (\pi -x))=1\) \(\implies - \sin x- \sin (\pi -x)=1\) \(\implies - \sin x -\sin x=1 \implies -2 \sin x=1 \implies \sin x=-\frac{1}{2}.\) \[sinx = -\sin (\frac{\pi}{6}) \implies sinx = \sin(\frac{- \pi}{6})\] Here the general solution is given by: \[x = n \pi + (-1)^n (\frac{- \pi}{6})\] Put the values of n as 0 , 1 , 2 etc etc..

OpenStudy (anonymous):

Put n = 0: \[x = \frac{- \pi}{6} \implies \frac{11 \pi}{6}\]

OpenStudy (anonymous):

Put n = 1: \[x = \pi + \frac{\pi}{6} \implies x = \frac{7 \pi}{6}\]